OpenStudy (anonymous):

5/7 divided by 1 4/7= find the estimate answer and then the exact answer

OpenStudy (anonymous):

5/7 approximate .7 14/7 approximate 2 so estimate = .7/2 = .35 5/7*7/14 = 5/14 which equals 0.357142857